# Quantum Corrections to Randall-Sundrum Model from JT Gravity

### v1

@@ I Introduction
The quest to unify the fundamental forces of nature and resolve the hierarchy problem, one of the most important problems in particle physics, has spurred the development of numerous higher-dimensional theoretical frameworks [DOMAIN-WALL, EXTRA1, WOS:000074694500008]. Among these, the Randall-Sundrum (RS) model, proposed by Lisa Randall and Raman Sundrum in 1999 [Large, alternative], stands as the first theory of warped extra dimensions, proposes a potential solution to the hierarchy problem, and has attracted attention in the academic community.
@@
The RS model exhibits a deep connection with the AdS/CFT correspondence [WOS:000081063400007], as the extra-dimensional evolution along the bulk maps directly to the renormalization group flow of a four-dimensional nearly conformal field theory (CFT) in the dual picture. A critical dynamical degree of freedom in the RS model is the radion, a massless mode associated with fluctuations of the radius of extra dimension, whose stabilization is described by the Goldberger-Wise (GW) mechanism [WOS:000084152000004]. Recent renewed interest in the RS model stems from its implications for early universe cosmology, including the generation of a stochastic gravitational wave background via supercooled phase transitions [WOS:000175277100051, WOS:001116716600002, WOS:000423800200001]. However, recent studies have indicated that it is a cooled phase transition rather than supercooled [WOS:001251818200003], further underscoring its relevance to both theoretical and observational physics.
@@
The JT gravity [JACKIW, twoSPACETIME] is a two-dimensional gravitational theory characterized by a simple action involving an Einstein-Hilbert term and a constant negative cosmological constant in the bulk as well as a Gibbons-Hawking-York term on the boundary. It has emerged as a powerful tool for describing the near-horizon dynamics of higher-dimensional near-extremal Reissner-Nordström (RN) black holes [Turiaci]. A defining insight is that the collective behavior of the near-horizon region of such a black hole, where the spacetime geometry reduces to AdS2, is effectively captured by a JT-like gravitational action, with the boundary dynamics governed by a Schwarzian action [Zhenbin]. The Schwarzian mode, arising from the spontaneous breaking of conformal symmetry in the AdS2 boundary, encodes crucial information about the thermodynamics of the underlying black hole, including its statistical entropy and free energy, which are correctly reproduced by the partition function of the Schwarzian action. For the Schwarzian action, we can perform a path integral, allowing us to consider quantum fluctuations in the near-horizon region of near-extremal black holes. This kind of quantum fluctuation stands out because thermal fluctuations are suppressed; we refer to it as an infrared quantum fluctuation. Recently, there have been papers discussing the corrections to various physical quantities caused by this type of quantum fluctuations [liu2024quantumcorrectionsholographicstrange, Brown:2024ajk, liu2025quantumcorrectedholographicwilsonloop, Emparan:2025sao, zheng, nian2025quantumcorrectionslowtemperaturefluidgravity, PandoZayas:2025snm, cremonini2025quantumcorrectionsetasjt]. In our paper, we investigate the quantum corrections to the RS model due to the quantum fluctuations of the near-extremal RN black brane.
@@ II Randall-Sundrum model
The Randall-Sundrum (RS) model posits a single extra dimension that is compactified on a circle, with the upper and lower halves of the circle identified to construct the $S^{1}/Z_{2}$ orbifold [Large]. Denoted by the coordinate $y$, this extra dimension possesses two fixed points located at $y=0$ and $y=\pi r_{c}$, where $r_{c}$ serves as the radius of the compactified extra dimension. Each fixed point is inhabited by a four-dimensional brane (referred to as a 3-brane), and these two 3-branes, spaced $\pi r_{c}$ apart, confine a five-dimensional (5D) bulk spacetime. The total action of the model is composed of three parts: one arising from gravity in the 5D bulk, and the other two encoding the physics localized on each of the two 3-branes, i.e.,
@@
where $g_{0}$ and $g_{\pi r_{c}}$ represent the determinants of the induced metrics of the 5D metric on the two branes, respectively. $\mathcal{L}$ represents the Lagrangian of the matter field on the brane, while $V$ is the energy density of the brane. The spacetime metric $G_{MN}$ is [Large]
@@
Note that the spacetime in between the two 3-branes is simply a slice of an AdS5 geometry. According to the AdS/CFT correspondence, the 5D bulk spacetime is dual to the 4D boundary conformal field theory [WOS:000081063400007]. The two branes correspond to the ultraviolet and the infrared cutoffs of the field theory. We use subscripts UV and IR to label the actions of the two branes located at $y=0$ and $y=\pi r_{c}$, respectively.
@@
The linearized Einstein equation in $y$ coordinate is [alternative]
@@
and $f_{0}=const$ for $m_{0}=0$. For $m_{n}\neq 0$, the general solution of the KK mode can be written as [alternative]
@@
where $N_{n}$ and $b_{n}$ are constants. Correspondingly, the KK mass spectrum is [alternative]
@@
Since $ke^{-k\pi r_{c}}\simTeV$, the KK gravitons in the RS model have masses of the TeV scale.
Up to this point, the radius of the extra dimension has been treated as a free parameter. However, this degree of freedom implies a massless scalar field in the effective theory, corresponding to fluctuations of the radius along the extra dimension. This massless scalar field would induce a fifth force in violation of the equivalence principle. Therefore, to preserve the viability of the Randall-Sundrum model, the scalar field has to obtain a mass, i.e., to be stabilized. A well-established approach to achieve this is the Goldberger-Wise mechanism [WOS:000084152000004]. Classically, a potential emerges from a bulk scalar field with interaction terms localized on the two 3-branes. This potential’s minimum can be tuned to yield $kr_{c}\sim 10$ without recourse to fine-tuning of parameters [WOS:000084152000004].
@@ III JT gravity and quantum correction to AdS2
Let us consider the Euclidean Jackiw-Teitelboim (JT) gravity action [JACKIW, twoSPACETIME]:
@@
where $R_{c}$ is the scalar curvature, and $K$ is the boundary extrinsic curvature. We first vary the action $I_{JT}$ with respect to the dilaton field $\phi$, leading to $R_{c}=-2$, which implies that the bulk is a patch of the 2d hyperbolic surface $ds^{2}=\frac{dT^{2}+dZ^{2}}{Z^{2}}$. Dynamics is now entirely dictated by the Gibbons-Hawking-York (GHY) boundary term on a boundary curve. This can be visualized on the hyperbolic upper half surface, where a region close to the actual boundary is cut out. We parametrize the boundary curve as $(T=F(s),Z(s))$, where $s$ is the proper time on the boundary. Fixing the boundary metric to $\frac{1}{\epsilon}$ imposes $Z(s)=\epsilon F^{\prime}(s)$. We can evaluate the extrinsic curvature [Zhenbin]:
@@ IV RS metric with quantum corrections
The quantum-corrected equation of motion is then given by the Schwinger-Dyson equation [Dyson, SCHWINGER]:
@@
In principle, we need to evaluate the path integral on (21) to determine $\langle A_{cor}\rangle$. Noting $\epsilon(x^{+})\epsilon(x^{-})=\frac{1}{2}(\epsilon(x^{+})\epsilon(x^{-})+\epsilon(x^{-})\epsilon(x^{+}))$, and considering that the one-point function of the Schwarzian mode vanishes, while the two-point function is given by [Sachdev-Ye-Kitaev, Zhenbin]
@@
If we include higher-order terms in the correction factor, we can obtain the mass spectrum by numerical solution. In order to solve the hierarchy problem, we require $1/L$ to be at the Planck scale, and $\pi r_{c}/L=36.84$ [Large]. In numerical calculations, we adopt the Planck scale, $10^{16}$ TeV, as the fundamental unit and set $L=1$. Perturbative calculations require $\beta/C$ to be less than 1, so we take a few values $0.1$, $0.9$, and $0.99$ for $\beta/C$. The limit $u\gg u_{T}$ means $\frac{L^{2}}{u_{0}\beta}e^{y/L}\ll\frac{2-Q^{2}}{6\pi}$, and we take $\frac{L^{2}}{u_{0}\beta}e^{\pi r_{c}/L}=0.1$, $0.05$, and $0.01$ accordingly. The numerical results are listed in the Tab. 1 through Tab. 6. Since we take the Planck scale as the fundamental unit and the eigenvalues $M_{n}^{2}$ are extremely small, we multiply them by $10^{32}$ in practice. Therefore, the unit of $M_{n}$’s numerical results should be $10^{-16}\,L^{-1}$, i.e., $\sim$ TeV.
@@ VI Goldberger-Wise mechanism
If we only consider the lowest order of the quantum correction factor, the effective mass does not vary with $y$, and we can use the well-known result to obtain [WOS:000084152000004]
@@
It is noted that a recent work [WOS:000970026700004] has also discussed the relationship between the RS model and the JT gravity, but there are differences between this work and our paper. In our paper, the JT gravity describes only the near-horizon geometry of black branes, not the braneworld model. We use JT gravity only as a tool to account for quantum fluctuations near the black brane horizons.

### v2

@@ I Introduction
The quest to unify the fundamental forces of nature and resolve the hierarchy problem, one of the most important problems in particle physics, has spurred the development of numerous higher-dimensional theoretical frameworks [23, 24, 1]. Among these, the Randall-Sundrum (RS) model, proposed by Lisa Randall and Raman Sundrum in 1999 [22, 21], stands as the first theory of warped extra dimensions, proposes a potential solution to the hierarchy problem, and has attracted attention in the academic community.
@@
The RS model exhibits a deep connection with the AdS/CFT correspondence [15], as the extra-dimensional evolution along the bulk maps directly to the renormalization group flow of a four-dimensional nearly conformal field theory (CFT) in the dual picture. A critical dynamical degree of freedom in the RS model is the radion, a massless mode associated with fluctuations of the radius of extra dimension, whose stabilization is described by the Goldberger-Wise (GW) mechanism [9]. Recent renewed interest in the RS model stems from its implications for early universe cosmology, including the generation of a stochastic gravitational wave background via supercooled phase transitions [4, 18, 3]. However, recent studies have indicated that it is a cooled phase transition rather than supercooled [19], further underscoring its relevance to both theoretical and observational physics.
@@
The JT gravity [11, 26] is a two-dimensional gravitational theory characterized by a simple action involving an Einstein-Hilbert term and a constant negative cosmological constant in the bulk as well as a Gibbons-Hawking-York term on the boundary. It has emerged as a powerful tool for describing the near-horizon dynamics of higher-dimensional near-extremal Reissner-Nordström (RN) black holes [10]. A defining insight is that the collective behavior of the near-horizon region of such a black hole, where the spacetime geometry reduces to AdS2, is effectively captured by a JT-like gravitational action, with the boundary dynamics governed by a Schwarzian action [16]. The Schwarzian mode, arising from the spontaneous breaking of conformal symmetry in the AdS2 boundary, encodes crucial information about the thermodynamics of the underlying black hole, including its statistical entropy and free energy, which are correctly reproduced by the partition function of the Schwarzian action. For the Schwarzian action, we can perform a path integral, allowing us to consider quantum fluctuations in the near-horizon region of near-extremal black holes. This kind of quantum fluctuation stands out because thermal fluctuations are suppressed; we refer to it as an infrared quantum fluctuation. Recently, there have been papers discussing the corrections to various physical quantities caused by this type of quantum fluctuations [13, 2, 14, 7, 12, 20, 27, 5]. In our paper, we investigate the quantum corrections to the RS model due to the quantum fluctuations of the near-extremal RN black brane.
@@ II Randall-Sundrum model
The Randall-Sundrum (RS) model posits a single extra dimension that is compactified on a circle, with the upper and lower halves of the circle identified to construct the $S^{1}/Z_{2}$ orbifold [22]. Denoted by the coordinate $y$, this extra dimension possesses two fixed points located at $y=0$ and $y=\pi r_{c}$, where $r_{c}$ serves as the radius of the compactified extra dimension. Each fixed point is inhabited by a four-dimensional brane (referred to as a 3-brane), and these two 3-branes, spaced $\pi r_{c}$ apart, confine a five-dimensional (5D) bulk spacetime. The total action of the model is composed of three parts: one arising from gravity in the 5D bulk, and the other two encoding the physics localized on each of the two 3-branes, i.e.,
@@
where $g_{0}$ and $g_{\pi r_{c}}$ represent the determinants of the induced metrics of the 5D metric on the two branes, respectively. $\mathcal{L}$ represents the Lagrangian of the matter field on the brane, while $V$ is the energy density of the brane. The spacetime metric $G_{MN}$ is [22]
@@
Note that the spacetime in between the two 3-branes is simply a slice of an AdS5 geometry. According to the AdS/CFT correspondence, the 5D bulk spacetime is dual to the 4D boundary conformal field theory [15]. The two branes correspond to the ultraviolet and the infrared cutoffs of the field theory. We use subscripts UV and IR to label the actions of the two branes located at $y=0$ and $y=\pi r_{c}$, respectively.
@@
The linearized Einstein equation in $y$ coordinate is [21]
@@
and $f_{0}=const$ for $m_{0}=0$. For $m_{n}\neq 0$, the general solution of the KK mode can be written as [21]
@@
where $N_{n}$ and $b_{n}$ are constants. Correspondingly, the KK mass spectrum is [21]
@@
The RS model assumes that the fundamental mass scale is the Planck scale. To solve the hierarchy problem, the parameter $k$ with mass dimension should also be on the Planck scale, and the effective mass scale on the 4D IR brane must be redshifted down to the TeV scale relative to the 5D fundamental scale. Consequently, $ke^{-k\pi r_{c}}$ is $\mathcal{O}(1\,\mathrm{TeV})$, and the KK gravitons in the RS model have masses of the TeV scale.
Up to this point, the radius of the extra dimension has been treated as a free parameter. However, this degree of freedom implies a massless scalar field in the effective theory, corresponding to fluctuations of the radius along the extra dimension. This massless scalar field would induce a fifth force in violation of the equivalence principle. Therefore, to preserve the viability of the Randall-Sundrum model, the scalar field has to obtain a mass, i.e., to be stabilized. A well-established approach to achieve this is the Goldberger-Wise mechanism [9]. Classically, a potential emerges from a bulk scalar field with interaction terms localized on the two 3-branes. This potential’s minimum can be tuned to yield $kr_{c}\sim 10$ without recourse to fine-tuning of parameters [9].
@@ III JT gravity and quantum correction to AdS2
Let us consider the Euclidean Jackiw-Teitelboim (JT) gravity action [11, 26]:
@@
where $R_{c}$ is the scalar curvature, and $K$ is the boundary extrinsic curvature. We first vary the action $I_{JT}$ with respect to the dilaton field $\phi$, leading to $R_{c}=-2$, which implies that the bulk is a patch of the 2d hyperbolic surface $ds^{2}=\frac{dT^{2}+dZ^{2}}{Z^{2}}$. Dynamics is now entirely dictated by the Gibbons-Hawking-York (GHY) boundary term on a boundary curve. This can be visualized on the hyperbolic upper half surface, where a region close to the actual boundary is cut out. We parametrize the boundary curve as $(T=F(s),Z(s))$, where $s$ is the proper time on the boundary. Fixing the boundary metric to $\frac{1}{\epsilon}$ imposes $Z(s)=\epsilon F^{\prime}(s)$. We can evaluate the extrinsic curvature [16]:
@@ IV RS metric with quantum corrections
The quantum-corrected equation of motion is then given by the Schwinger-Dyson equation [6, 25]:
@@
In principle, we need to evaluate the path integral on (21) to determine $\langle A_{cor}\rangle$. Noting $\epsilon(x^{+})\epsilon(x^{-})=\frac{1}{2}(\epsilon(x^{+})\epsilon(x^{-})+\epsilon(x^{-})\epsilon(x^{+}))$, and considering that the one-point function of the Schwarzian mode vanishes, while the two-point function is given by [17, 16]
@@
If we include higher-order terms in the correction factor, we can obtain the mass spectrum by numerical solution. In order to solve the hierarchy problem, we require $1/L$ to be at the Planck scale, and $\pi r_{c}/L=36.84$ [22]. In numerical calculations, we adopt the Planck scale, $10^{16}$ TeV, as the fundamental unit and set $L=1$. Perturbative calculations require $\beta/C$ to be less than 1, so we take a few values $0.1$, $0.9$, and $0.99$ for $\beta/C$. The limit $u\gg u_{T}$ means $\frac{L^{2}}{u_{0}\beta}e^{y/L}\ll\frac{2-Q^{2}}{6\pi}$, and we take $\frac{L^{2}}{u_{0}\beta}e^{\pi r_{c}/L}=0.1$, $0.05$, and $0.01$ accordingly. The numerical results are listed in the Tab. 1 through Tab. 6. Since we take the Planck scale as the fundamental unit and the eigenvalues $M_{n}^{2}$ are extremely small, we multiply them by $10^{32}$ in practice. Therefore, the unit of $M_{n}$’s numerical results should be $10^{-16}\,L^{-1}$. Since $L^{-1}$ is on the Planck scale, the unit of numerical results in the table is TeV.
@@ VI Goldberger-Wise mechanism
If we only consider the lowest order of the quantum correction factor, the effective mass does not vary with $y$, and we can use the well-known result to obtain [9]
@@
It is noted that a recent work [8] has also discussed the relationship between the RS model and the JT gravity, but there are differences between this work and our paper. In our paper, the JT gravity describes only the near-horizon geometry of black branes, not the braneworld model. We use JT gravity only as a tool to account for quantum fluctuations near the black brane horizons.